Two fair cubes are rolled. The random variable X represents the difference between the values of the two cubes.

 

a)     Find the mean of this probability distribution. (i.e. Find E[X] )

b)     Find the variance and standard deviation of this probability distribution.

(i.e. Find V[X] and SD[X])

The random variables A and B are defined as follows:

A = X-10 and B = [(1/2)X]-5

c) Show that E[A] and E[B].

d) Find V[A] and V[B].

e) Arnold and Brian play a game using two fair cubes. The cubes are rolled, and Arnold records his score using the random variable A and Brian uses the random variable B. They repeat this for a large number of times and compare their scores. Comment on any likely differences or similarities of their scores.




1
Expert's answer
2021-11-22T15:12:02-0500

The random variable "X" may take on the following difference between the values of the two cubes.

"x=0,1,2,3,4,5" and its probability distribution is,

"x" 0 1 2 3 4 5

"p(x)" 1/6 5/18 2/9 1/6 1/9 1/18


a.

The mean of the probability distribution above is,

"E(x)=\sum x*p(x)=(0*1/6)+(1*5/18)+(2*2/9)+(3*1/6)+(4*1/9)+(5*1/18)=70/36= 1.944444"


b.

To find the variance of the random variable "X," we need to find the value of "E(X^2)" given as,

"E(x^2)=\sum x^2*p(x)=(1*5/18)+(4*2/9)+(9*1/6)+(16*1/9)+(25*1/18)= 5.833333"

We now use the formula below to find the variance,

"V(X)=E(x^2)-(E(X))^2=5.833333-(1.944444)^2=2.052471"

The standard deviation is given as,

"SD(X)=\sqrt{V(X)}=\sqrt{2.052471}= 1.432645"

Therefore the variance and the standard deviation of "X" are 5.833333 and 1.432645 respectively.

c.

Given that

A = X-10 and B = [(1/2)X]-5 then,

"E(A)=E(X-10)=E(X)-10=1.944444-10= -8.055556" and

"E(B)=E(1\/2*X-5)=1\/2*E(X)-5=(1\/2*1.944444)-5= -4.027778"

Therefore, "E(A)" and "E(B)" are -8.055556 and -4.027778 respectively.

d.

Given A = X-10 and B = [(1/2)X]-5 then,

"V(A)=V(X-10)=V(X)-V(10)=V(X)=2.052471" since the variance of a constant is zero.

"V(B)=V(1\/2*X-5)=(1\/2)^2V(X)-V(5)=1\/4*2.052471-0=1\/4*2.052471= 0.5131178"

Therefore, "V(A)" and "V(B)" are 2.052471 and 0.5131178 respectively.

e.

When Arnold records his score using the random variable A and Brian uses the random variable B, then the random variable "A\sim Binom(n,p)" where "n" is the number of times Arnold rolls the pair of fair cubes and "p" is the probability of success at each of the values that the random variable "A" can take. Also, the random variable "B\sim Binom(n,p)" where "n" is the number of times Brian rolls the pair of fair cubes and "p" is the probability of success at each of the values that the random variable "B" can take. Therefore, the probability distribution for rolling fair cubes for both will have the same Binomial distribution with parameters "n" and "p."